PREFACE

In succession to former conferences in mathematics sponsored by Josai University, the First International Symposium on Differential Geometry took place at the Conference Hall of Mizuta Memorial Library, Josai University from February 22 to February 24, 2001. The scientific programme consists of sixteen lectures listed below.

The present volume contains sixteen contributions from an impressive group of geometers across a wide range of topics in Riemannian geometry, theory of submanifolds, harmonic maps and integrable systems. We wish that these proceedings would provide the readers with a good view of the present status of differential geometry in many aspects.

It is our pleasant duty to thank Josai University for the financial support. We would also like to thank all the speakers for their enthusiastic participation and their cooperation in writing up their talks. Thanks are also due to the university staff for their various support.

Qing-Ming Cheng

Josai University